# Birgir Hrafnkelsson Morris Riedel Egill Skúlason

### v1

@@ 5.5 Cataloging saddles
This diagnostic protocol demonstrates that the Cartesian coordinate single-ended saddle searches, efficiently identify the proximal barriers that govern basin-to-basin transitions in complex systems, while internal coordinate systems may struggle. Validation of results in a chemical context with efficient tools such as the MMF-NEB must be undertaken, as common point measures of success (single negative eigenvalue, convergence, “barrier” from initial) may be misleading as shown here. In the twenty-first century, visual analysis is infeasible, even for the 500 systems described in this thesis (Section 5.3). Typical line plot comparisons and standard errors are dangerously oversimplified for the physical chemistry context, and we address this in the next chapter with Bayesian hierarchical models applied to modalities of dimer searches.
@@ 6 Dimer rotations and Hierarchical Bayesian models
This chapter is based on [95]
@@ 6.2.1 Computational Effort
We modeled the number of calls for the Potential Energy Surface using a negative binomial distribution, as specified in Eq. 56.
@@ 6.2.2 Wall time estimates
As time is a continuous and strictly positive quantity, we employed a Gamma generalized linear mixed model, the appropriate distributional choice for such data. The model specification is shown in Eq. 57.
@@ 6.2.3 Convergence Success
We analyzed the probability of a successful search using a Bernoulli logistic regression model, detailed in Eq. 58.
@@ 7 Data efficiency for Gaussian Processes
In this chapter we discuss practical amelioration techniques for the theoretical scaling of the A collection of random variables, any finite number of which have a joint multivariate normal (Gaussian) distribution; defines a distribution over functions. (section 2.3.1). Starting with lossless methods, both implemented and not implemented in the gpr_optim we move to lossy approximations, before setting the stage for the data-efficient methodology developed in the final part of the thesis.
To recap, each training geometry provides both energy and force information, leading to a total of $3N+1$ targets per configuration for a molecule with $N$ atoms. For $M$ training geometries, the classic formulation requires constructing and inverting the full covariance matrix $\mathbf{K}$ of size $(M(3N+1))\times(M(3N+1))$. This results in a memory cost scaling as $\mathcal{O}(M^{2}N^{2})$ and computational cost (for matrix inversion or Cholesky decomposition) scaling as $\mathcal{O}(M^{3}N^{3})$. Such cubic scaling quickly becomes infeasible as either $M$ or $N$ increases.
@@ 7.1 Quicker inversions through reshaping
However, in practice; and, in this work, the optimization routine does not require the explicit inversion of the full kernel matrix. Instead, the marginal likelihood and its gradient with respect to kernel hyperparameters are evaluated using efficient matrix-vector products and decompositions. This is achieved by assembling a rectangular matrix $\mathbf{R}$ of size $M(3N+1)\times(3N+1)$, in which each row corresponds to a target from the training set, and each column corresponds to a specific observable (energy or a force component).
In linear algebra, reordering or reshaping a matrix is mathematically a no-op with respect to fundamental operations such as inversion, decomposition, and linear solves [105]. Any permutation or block reorganization of a matrix $\mathbf{K}$ can be written as
@@
where $\mathbf{P}$ is a permutation matrix (orthogonal, $\mathbf{P}^{T}=\mathbf{P}^{-1}$). For any vector $\mathbf{y}$,
@@
Consider a “reshaped” or block-organized version of $\mathbf{K}$, e.g., storing targets grouped by configuration or observable. The Cholesky decomposition (used in SCG and GP marginal likelihood) and matrix-vector solves are unaffected:
@@
Thus, the storage and computational cost of this approach scale as $\mathcal{O}(MN^{2})$, which is linear in the number of training geometries and quadratic in the number of atoms. This reduction is achieved without any loss of information or accuracy, as all targets are still included in the optimization; the difference is purely the result of efficient organization and evaluation of the required matrix operations. Scaling from a softare design perspective is known to unlock linear scaling in computational chemistry [106].
@@
Figure 23 contrasts the practical scaling of the block matrix used in this implementation (solid lines and dots) with the theoretical full kernel scaling (dashed lines). The block matrix size is plotted as both lines and points, while the theoretical scaling is shown as dashed lines for each molecule size. The improved scaling of the block approach enables GPR models to be trained on much larger systems and datasets.
Figure 23: Number of elements in kernel matrices for Gaussian Process training with energy and force data. Solid lines with points show the practical block matrix ($M(3N+1)\times(3N+1)$), while dotted lines show the theoretical full kernel ($(M(3N+1))^{2}$). The number of elements is expressed as powers of 10 versus the number of training geometries, for several molecule sizes. For example, with $N=18$ atoms and $M=75$ geometries, the block matrix contains 226,875226,\!875 elements ($10^{5.4}$), compared to 17,015,62517,\!015,\!625 elements ($10^{7.2}$) for the full kernel.
@@ 7.2 Rank one covariance updates for new data
When hyperparameters $\boldsymbol{\theta}$ are held fixed, adding a new observation $(\mathbf{x}_{n+1},y_{n+1})$ to an existing dataset of size $n$ constitutes a rank-one update to the covariance matrix. The inverse of this new matrix can be computed efficiently from the old inverse using the Woodbury matrix identity, or more stably, its Cholesky factor can be updated in $O(n^{2})$ time. This avoids the full $O(n^{3})$ cost of re-decomposition. The update is algebraically exact and information-lossless.
A conceptual Cholesky update proceeds as follows: given the old Cholesky factor $\mathbf{L}_{n}$ of the $n\times n$ matrix $\mathbf{K}_{n}$, the new factor $\mathbf{L}_{n+1}$ has the block structure:
$\mathbf{L}_{n+1}=\begin{pmatrix}\mathbf{L}_{n}&\mathbf{0}\ \mathbf{l}^{T}&l_{d}\end{pmatrix}$ (65)
where $\mathbf{l}$ is found by solving $\mathbf{L}_{n}\mathbf{l}=\mathbf{k}$ and $l_{d}=\sqrt{k_{dd}-\mathbf{l}^{T}\mathbf{l}}$. In the gpr_optim this approach is not pursued because the formulation of a rank one update proceedure for adding the training covariance is overshadowed by the subsequent re-optimization of the hyperparameters, where the entire matrix is inverted repeatedly.
The search for optimal hyperparameters, $\boldsymbol{\theta}^{*}$, requires maximizing the marginal likelihood:
$\boldsymbol{\theta}^{*}=\arg\max_{\boldsymbol{\theta}}\left(-\frac{1}{2}\mathbf{y}^{T}\mathbf{K}{\boldsymbol{\theta}}^{-1}\mathbf{y}-\frac{1}{2}\log\det(\mathbf{K}{\boldsymbol{\theta}})\right)$ (66)
Crucially, the covariance matrix $\mathbf{K}_{\boldsymbol{\theta}}$ is a function of the hyperparameters. Each candidate set of hyperparameters $\boldsymbol{\theta}_{i}$ in the search defines an entirely new matrix. The incremental update is rendered useless because the base matrix $\mathbf{K}_{n}$ is not fixed; it is constantly being redefined.
@@ 7.3 Pruning over data
The coupling between data support (Eq. 73) and inference (Eq. 71, 72) admits no free lunch.
@@ 7.4 Variance and accuracy
We draw $N_{\text{trials}}=60$ candidate hyperparameters uniformly in log-space and select the maximum.
@@
Figures 25 and 27 present the empirical comparison on the Himmelblau function (Eq. 75). $T=28$ observations are taken along a random walk initialized at $\mathbf{x}_{0}=(-3.0,3.0)$, a point deep in a high-valued basin. A dense probe grid of $45\times 45$ points covers the domain $6,6]^{2}$.
@@
Taken together, these results demonstrate a clear pathology in naively re-optimizing GPDimer hyperparameters within a sequential learning process, particularly when using derivative information. While re-optimization can improve global poinit-wise accuracy metrics, it does so by sacrificing the integrity of the model’s variance estimates. The resulting model becomes volatile and severely overconfident, producing miscalibrated uncertainty predictions that cannot be trusted for decision-making. This reinforces the need for caution when interpreting predictive variance from models whose hyperparameters are continually adapted on growing, locally-clustered datasets.
@@ 7.5.1 Hyperparameter trajectories for the GPDimer
The hyperparameters in GPDimer runs are seen to stabilize after a modest amount of data 191919though these are not globally optimal like thos from optuna [109]. To this end, we employ a local, gradient-based SCG optimizer for the hyperparameters (Alg. 4), warm-started from the converged values of the previous optimization step. This avoids the cost of a global search and leverages the fact that the Potential Energy Surface topology evolves smoothly.
@@ 7.5.3 Hyperparameter Oscillation Detection
7.6 Summary
The scaling of Gaussian Process methods relies on lossless approximations rooted in software design, hardware acceleration, or lossy approximations. The cost of inversion falls disproportionately on the hyperparameter optimization, and so a key insight developed here in this thesis involves pruning for this step. In the final section, we complete the framework for wall-time-efficient Gaussian Process with the introduction of a distance measure to select points, needed for both pruning and early stopping.
@@ 8 Optimal Transport Gaussian Process
The Gaussian Process, defined in terms of finite realizations of multivariate normal distributions, expressed as a “function”, or a series of $x,y$ pairs reshaped to provide familiarity with 3D matrices; thus suffers greatly from anthropomorphism of the constituent equations. The physical meaning ascribed to the hyperparameters are unjustified, and models will only be guaranteed to interpolate in the noise free regime 202020also known to be unstable [33]. The repeated re-optimization also doesn’t preserve global accuracy as shown earlier. This distinction becomes decisive in actively learned saddle-point searches, which produce correlated trajectories of geometries which are atypical from most of the energy surface geometrically. Under such data, the MLL surface can be shallow in directions like the signal variance, encouraging variance blow-up that flattens the mean and inflates predictive uncertainty, which contributes greately to the destabilization of the GPDimer. From a data efficiency perspective, what was covered in section 7 has even greater significance in terms of reliability. However, we still require a reasonable measure of distance to complete algorithm developed in this chapter to make good on wall time performance, which we summarize in Fig. 31.

### v2

@@ 5.6 Conclusions
Validation of results in a chemical context with efficient tools such as the MMF-NEB must be undertaken, as common point measures of success (single negative eigenvalue, convergence, “barrier” from initial) may be misleading as shown in the case study. In the twenty-first century, visual analysis is infeasible, even for the 500 systems described in this thesis (Section 5.3). Typical line plot comparisons and standard errors are dangerously oversimplified for the physical chemistry context, and we address this in the next chapter with Bayesian hierarchical models applied to modalities of dimer searches.
@@ 6 Dimer rotations and Hierarchical Bayesian models
This chapter is based on [138] with applications in [139]
@@ 6.2.1 Computational Effort
We modeled the number of calls for the Potential Energy Surface using a negative binomial distribution, as specified in Eq. 57. This choice is motivated by the physical constraints of the metric: Potential Energy Surface calls are discrete, non-negative integers. Standard linear models assume normality, which allows for non-integers and negative values. Furthermore, algorithmic performance data typically exhibits overdispersion, where the variance exceeds the mean—due to the heterogeneous nature of chemical landscapes. The Negative Binomial distribution explicitly models this overdispersion, preventing the underestimation of standard errors common in Poisson or Gaussian approximations.
@@ 6.2.2 Wall time estimates
Wall time is a strictly positive, continuous variable. The distribution of runtimes is often right-skewed, featuring a long tail of difficult systems while the majority converge quickly. A standard Gaussian model would imply a symmetric distribution, potentially assigning non-zero probability to physically impossible negative times. We therefore employed a Gamma generalized linear mixed model, which is naturally defined for $x>0$ and flexibly accommodates the skewness inherent in computational timing data. The model specification is shown in Eq. 58.
@@ 6.2.3 Convergence Success
We analyzed the probability of a successful search using a Bernoulli logistic regression model, detailed in Eq. 59. As convergence is a binary event ($y\in\{0,1\}$), the search either isolates a saddle point or it does not, the Bernoulli distribution is the canonical probability model for such dichotomous outcomes.
@@ 6.3 Revisiting rotation removal
6.4 Time-to-Solution as a Function of Effort
While independent analyses of computational effort (PES calls) and wall time provide valuable baseline metrics, they do not fully capture the dynamic relationship between the two. In a practical optimization scenario, the “cost” of a method is not just how many steps it takes, but how the time-per-step scales as the optimization trajectory lengthens.
To investigate this, we modeled the total time-to-solution ($T$) as a continuous function of the computational effort ($N_{\text{calls}}$), conditioned on the optimization method. We employed a Bayesian Hierarchical Gamma Spline model (Eq. 66) to capture the potentially non-linear scaling behaviors while accounting for the strict positivity and skewness of runtime data.
$T_{ij}$ $\sim\text{Gamma}(\mu_{ij},\alpha)$ (66)
$\log(\mu_{ij})$ $=\beta_{0}+\beta_{\text{method}}+f_{\text{method}}(\log(N_{\text{calls},ij}))+u_{j}$
Here, $f_{\text{method}}$ represents a thin-plate regression spline smooth term allowed to vary by method. This approach enables us to separate the fixed overhead of a method (intercept differences) from its scaling behavior (the shape of the spline). We also incorporated random intercepts $u_{j}$ for each system to control for the intrinsic expense of different chemical environments.
6.4.1 Extrapolation and Data Efficiency
A challenge in benchmarking machine-learning enhanced methods (like GPDimer and OTGPD, introduced in Chapter 8) against classical baselines (Dimer) is the disparity in convergence speed. The ML methods often converge with significantly fewer calls, leaving no observed data in the “high-effort” regime populated by the classical method.
Figure 24 visualizes this relationship. The solid lines represent the interpolated predictions within the observed data range for each method, while the dashed lines indicate extrapolation.
Figure 24: Modelled Time-to-Solution as a Function of Computational Effort. Posterior predictions from a Gamma spline model. The solid curves indicate the interpolation range where data was observed; dashed curves indicate extrapolation. The vertical dashed lines mark the maximum observed PES calls for the GP-based methods.
The plot reveals a distinct crossover. For extremely short runs (low PES calls), the ML-based methods incur a higher time cost due to the overhead of Gaussian Process regression. However, as the optimization difficulty increases (moving right on the x-axis), this overhead is amortized. It is also worth noting that the absolute time per call here is very low, since the Hartree-Fock calculations considered in this thesis do not take much time per configuration. Systems with more complex potential energy calculations will perform far better under the GP as the Dimer methods take an order of magnitude more calls.
6.4.2 Quantitative Comparison
To rigorously quantify this trade-off, we compared the expected time-to-solution at distinct regimes of computational effort, the overall results of which are shown in Table 4.
First, we evaluated the methods at a fixed budget of 30 PES calls, a typical early-phase regime for GP methods. At this point, the classical Dimer method is slightly faster (11.5 min) compared to GPDimer (12.8 min), reflecting the lower per-step cost of the classical algorithm. The OTGPD method, however, effectively mitigates this overhead (11.3 min), performing on par with the classical baseline even in this low-data regime.
However, comparing methods at a fixed x-value is artificial if the methods require vastly different amounts of effort to solve the same problem. A more physically meaningful comparison evaluates each method at its median convergence effort which approximates the typical “price” to solve a system.
Table 4: Results of modelled time to solution as a function of computational effort.
Method Median PES Calls Expected Time (min) 95% CrI
OTGPD 28 9.0 [8.6, 9.5]
GPDimer 30 12.8 [12.2, 13.5]
Dimer 254 20.5 [19.2, 21.8]
At their respective operating points, the efficiency gain is stark. The classical Dimer requires typically 254 calls, resulting in an expected runtime of 20.5 minutes. The OTGPD method achieves solution with only 28 calls on average, dropping the total wall time to 9.0 minutes. This represents a factor of 2.2x speedup in real-world wall time, confirming that the dramatic reduction in force calls provided by the GP surrogate translates directly to time savings even for trivial energy surfaces, despite the added computational overhead of the machine learning model.
6.5 Conclusions
This chapter established a rigorous statistical framework for evaluating computational performance, moving the discourse from anecdotal benchmarking to hierarchical Bayesian inference. By employing generalized linear mixed models with appropriate response distributions—Negative Binomial for effort, Gamma for wall time, and Bernoulli for convergence—we quantified the distinct trade-offs inherent in algorithmic choices while explicitly accounting for system-specific variability.
The analysis of the dimer method’s rotation phase yielded decisive results. We demonstrated that the Conjugate Gradient optimizer offers superior robustness compared to L-BFGS, contradicting default choices in established software packages. Furthermore, we showed that the specific implementation of rotation removal within the EON suite incurred a significant computational penalty without providing a credible improvement in convergence success.
However, the inefficiency of one specific implementation does not negate the underlying physical principle. As discussed in Section 6.3, the explicit removal of external invariances functions as a critical stability mechanism for Gaussian Process surrogates [139], preventing unphysical extrapolations in the active learning loop. This distinction underscores a broader theme in computational method development: the utility of a physical constraint depends heavily on its numerical integration within the optimization framework.
Using the same framework, modeling the time-to-solution landscape (Section 6.4) resolves the “overhead vs. efficiency” debate for machine-learning enhanced optimization. While surrogate-based methods like GPDimer and OTGPD incur a higher computational cost per step for cheap potentials evident in the crossover observed in the Gamma spline analysis, this overhead is effectively amortized over the trajectory. Our analysis confirms that the On-The-Fly Gaussian Process Dimer (OTGPD) method reduces the median computational effort so drastically (from ~254 to ~28 calls) that it yields a net reduction in total wall time of approximately 56% compared to the classical baseline. This result statistically validates the core premise of this thesis: that data efficiency translates directly to wall-clock acceleration, provided the surrogate model is strictly stabilized.
With a validated statistical methodology to measure success and a refined understanding of how to handle physical invariances, we now address the primary bottleneck in machine-learning-guided optimization, that of total wall time. The following chapter applies these insights to the challenge of data efficiency, developing strategies to scale Gaussian Process surrogates to chemically relevant system sizes.
@@ 7 Data efficiency for Gaussian Processes
A collection of random variables, any finite number of which have a joint multivariate normal (Gaussian) distribution; defines a distribution over functions. (Section 2.4.1) surrogates promise analytic posteriors and gradients, yet their naive scaling, $O(M^{3}N^{3})$ in time and $O(M^{2}N^{2})$ in memory for $M$ geometries and $N$ atoms, blocks routine use. This chapter develops a narrative from lossless algebraic reorganizations that preserve fidelity, to lossy approximations that trade information for wall-time. Along the way, we confront pathologies that emerge in active learning loops and establish safeguards that let the later chapters stand on firm computational ground.
@@ 7.1 Quicker inversions through reshaping
In this work, the optimization routine circumvents the explicit inversion of the full kernel matrix. Instead, the log MLL and its gradient with respect to kernel hyperparameters are evaluated using efficient matrix-vector products and decompositions. This is realized by assembling a rectangular matrix $\mathbf{R}$ of size $M(3N+1)\times(3N+1)$, where rows correspond to training targets and columns to specific observables.
Reordering or reshaping a matrix is algebraically a no-op regarding fundamental operations such as inversion and decomposition [149]. Any permutation or block reorganization of a matrix $\mathbf{K}$ can be written as
@@
where $\mathbf{P}$ is an orthogonal permutation matrix ($\mathbf{P}^{T}=\mathbf{P}^{-1}$). For any vector $\mathbf{y}$,
@@
Consider a “reshaped” or block-organized version of $\mathbf{K}$, e.g., storing targets grouped by configuration or observable. The Cholesky decomposition (used in SCG and Gaussian Process marginal likelihood) and matrix-vector solves are unaffected:
@@
Thus, the storage and computational cost of this approach scale as $\mathcal{O}(MN^{2})$, which is linear in the number of training geometries and quadratic in the number of atoms. This reduction is achieved without any loss of information or accuracy, as all targets are still included in the optimization; the difference is purely the result of efficient organization and evaluation of the required matrix operations. Scaling from a software design perspective is known to unlock linear scaling in computational chemistry [150].
@@
Figure 25 illustrates the practical scaling of the block matrix (solid lines) versus the theoretical full kernel (dashed lines). For a system with $N=18$ atoms and $M=75$ geometries, the full kernel requires $1.7\times 10^{7}$ elements ($10^{7.2}$), whereas the block matrix requires only $2.2\times 10^{5}$ elements ($10^{5.4}$). This efficiency gain enables GPR model training on significantly larger systems.
Figure 25: Kernel matrix element scaling for Gaussian Process training with energy and force data. Solid lines (with points) denote the practical block matrix ($M(3N+1)\times(3N+1)$), while dotted lines denote the theoretical full kernel ($(M(3N+1))^{2}$). Values are log-scale versus the number of training geometries. For $N=18$ atoms and $M=75$ geometries, the block matrix is orders of magnitude smaller than the full kernel formulation.
@@ 7.2 Rank one covariance updates for new data
With fixed hyperparameters, one new observation permits an $O(n^{2})$ Cholesky update rather than a full $O(n^{3})$ refactor:
$\mathbf{L}_{n+1}=\begin{pmatrix}\mathbf{L}_{n}&\mathbf{0}\\ \mathbf{l}^{T}&l_{d}\end{pmatrix},\quad\mathbf{L}_{n}\mathbf{l}=\mathbf{k},\quad l_{d}=\sqrt{k_{dd}-\mathbf{l}^{T}\mathbf{l}}.$ (67)
In the gpr_optim this approach is not pursued because the formulation of a rank one update procedure for adding the training covariance is overshadowed by the subsequent re-optimization of the hyperparameters, where the entire matrix is inverted repeatedly.
The search for optimal hyperparameters, $\boldsymbol{\theta}^{*}$, requires maximizing the marginal likelihood, but the covariance matrix $\mathbf{K}_{\boldsymbol{\theta}}$ is a function of the hyperparameters. Each candidate set of hyperparameters $\boldsymbol{\theta}_{i}$ in the search defines an entirely new matrix. The incremental update is rendered useless because the base matrix $\mathbf{K}_{n}$ is not fixed; it is constantly being redefined.
@@ 7.3 Pruning over data
The coupling between data support (Eq. 74) and inference (Eq. 72, 73) admits no free lunch.
@@ 7.4 Variance and accuracy
We draw $N_{\text{trials}}=60$ candidate hyperparameters uniformly in log-space and select the maximum. Figure 28 illustrates the resulting trajectories.
@@
Figures 27 and 29 present the empirical comparison on the Himmelblau function (Eq. 76). $T=28$ observations are taken along a random walk initialized at $\mathbf{x}_{0}=(-3.0,3.0)$, a point deep in a high-valued basin. A dense probe grid of $45\times 45$ points covers the domain $6,6]^{2}$.
@@
Taken together, these results demonstrate a clear pathology in naively re-optimizing GPDimer hyperparameters within a sequential learning process, particularly when using derivative information. While re-optimization can improve global point-wise accuracy metrics, it does so by sacrificing the integrity of the model’s variance estimates. The resulting model becomes volatile and severely overconfident, producing miscalibrated uncertainty predictions that cannot be trusted for decision-making. This reinforces the need for caution when interpreting predictive variance from models whose hyperparameters are continually adapted on growing, locally-clustered datasets.
@@ 7.5.1 Hyperparameter trajectories for the GPDimer
The hyperparameters in GPDimer runs are seen to stabilize after a modest amount of data 232323though these are not globally optimal like those from optuna [153]. To this end, we employ a local, gradient-based SCG optimizer for the hyperparameters (Alg. 4), warm-started from the converged values of the previous optimization step. This avoids the cost of a global search and leverages the fact that the Potential Energy Surface topology evolves smoothly.
@@ 7.5.3 Hyperparameter Oscillation Detection
7.6 Conclusions
This chapter demonstrated that scaling Gaussian Process methodologies for high-performance chemistry requires a synthesis of algebraic restructuring and careful statistical approximation. We established that matrix reshaping into block-diagonal forms reduces memory complexity to linear scaling without sacrificing mathematical fidelity. However, the wall-time cost of inverting these matrices during iterative hyperparameter optimization remains the primary computational bottleneck.
Our investigation into naive data pruning revealed severe pathologies; coupling the data support set directly to the inference engine creates feedback loops that drive optimization trajectories toward erroneous local minima. Furthermore, aggressive hyperparameter re-optimization on dynamic local subsets sacrifices variance calibration for marginal gains in point-wise accuracy, rendering the model overconfident and unsuitable for autonomous decision-making.
To mitigate these risks while capturing the immense speedups offered by sparse kernels—exceeding $7000\times$ for larger systems we introduced essential stability mechanisms. The adaptive signal variance barrier and Hyperparameter Oscillation Detection (HOD) prevent the unphysical divergence of the model during the volatile early stages of learning. These safeguards permit the use of reduced datasets for the hyperparameter optimization, yet the specific criterion for which data to retain remains critical. Simple Euclidean metrics fail to capture chemical similarity effectively across diverse landscapes. Consequently, the final component of this framework demands a rigorous, chemically transferable metric for point selection. Chapter 8 addresses this need through the introduction of the OTGPD, replacing geometric distance with Optimal Transport metrics to guide both pruning and convergence.
@@ 8 Optimal Transport Gaussian Process
The Gaussian Process, defined in terms of finite realizations of multivariate normal distributions, expressed as a “function”, or a series of $x,y$ pairs reshaped to provide familiarity with 3D matrices; thus suffers greatly from anthropomorphism of the constituent equations. The physical meaning ascribed to the hyperparameters are unjustified, and models will only be guaranteed to interpolate in the noise free regime 242424also known to be unstable [45]. The repeated re-optimization also doesn’t preserve global accuracy as shown earlier. This distinction becomes decisive in actively learned saddle-point searches, which produce correlated trajectories of geometries which are atypical from most of the energy surface geometrically. Under such data, the MLL surface can be shallow in directions like the signal variance, encouraging variance blow-up that flattens the mean and inflates predictive uncertainty, which contributes significantly to the destabilization of the GPDimer. From a data efficiency perspective, what was covered in section 7 has even greater significance in terms of reliability. However, we still require a reasonable measure of distance to complete algorithm developed in this chapter to make good on wall time performance, which we summarize in Fig. 33.